What is the hydronium ion concentration of 1g of acetic acid in 100 ml of water?


1
Expert's answer
2022-03-16T18:01:04-0400

Solution:

acetic acid = CH3COOH

hydronium ion = H3O+


The molar mass of CH3COOH is 60.052 g/mol

Therefore,

(1 g CH3COOH) × (1 mol CH3COOH / 60.052 g CH3COOH) = 0.01665 mol CH3COOH


Liters of CH3COOH solution = (100 mL soln) × (1 L / 1000 mL) = 0.1 L


Molarity of CH3COOH = Moles of CH3COOH / Liters of CH3COOH solution

Therefore,

Molarity of CH3COOH = (0.01665 mol) / (0.1 L) = 0.1665 mol/L = 0.1665 M


Acetic acid (CH3COOH) is a weak acid because it partially dissociates in an aqueous solution or it does not ionize completely in solution to produce hydronium ions (H3O+).

Equation representing the dissociation of CH3COOH:

CH3COOH(aq) + H2O(l) ⇋ H3O+(aq) + CH3COO(aq)

ICE Table:



According to the ICE Table:

[H3O+] = [CH3COO] = X

[CH3COOH] = 0.1665 − X ≈ 0.1665 M (assume that 0.1665 >> X)


The Ka expression for the dissociation of CH3COOH:


Ka = 1.75×10−5 (for CH3COOH)

Therefore,

1.75×10−5 = [X] × [X] / [0.1665]

2.914×10−6 = X2

X = (2.914×10−6)1/2 = 0.0017

[H3O+] = X = 0.0017 M = 1.7×10−3 M

[H3O+] = 1.7×10−3 M


Answer: The hydronium ion (H3O+) concentration is 1.7×10−3 M